# Accessible, At-Home Detection of Parkinson’s Disease via Multi-task Video Analysis

### v3

@@ Abstract
Limited accessibility to neurological care leads to underdiagnosed Parkinson’s Disease (PD), preventing early intervention. Existing AI-based PD detection methods primarily focus on unimodal analysis of motor or speech tasks, overlooking the multifaceted nature of the disease. To address this, we introduce a large-scale, multi-task video dataset consisting of $1102$ sessions (each containing videos of finger tapping, facial expression, and speech tasks captured via webcam) from $845$ participants ($272$ with PD). We propose a novel Uncertainty-calibrated Fusion Network (UFNet) that leverages this multimodal data to enhance diagnostic accuracy. UFNet employs independent task-specific networks, trained with Monte Carlo Dropout for uncertainty quantification, followed by self-attended fusion of features, with attention weights dynamically adjusted based on task-specific uncertainties. To ensure patient-centered evaluation, the participants were randomly split into three sets: $60\%$ for training, $20\%$ for model selection, and $20\%$ for final performance evaluation. UFNet significantly outperformed single-task models in terms of accuracy, area under the ROC curve (AUROC), and sensitivity while maintaining non-inferior specificity. Withholding uncertain predictions further boosted the performance, achieving $88.0\pm 0.3\%$ accuracy, $93.0\pm 0.2\%$ AUROC, $79.3\pm 0.9\%$ sensitivity, and $92.6\pm 0.3\%$ specificity, at the expense of not being able to predict for $2.3\pm 0.3\%$ data ($\pm$ denotes 95% confidence interval). Further analysis suggests that the trained model does not exhibit any detectable bias across sex and ethnic subgroups and is most effective for individuals aged between 50 and 80. Requiring only a webcam and microphone, our approach facilitates accessible home-based PD screening, especially in regions with limited healthcare resources.
@@ Introduction
Detecting PD is particularly challenging due to the variability of individual symptoms. For example, while vocal impairment is a common PD symptom (Ho et al. 1999), patients may also exhibit PD through other modalities, such as facial expression (e.g., hypomimia (Gunnery et al. 2016)) or motor function (e.g., bradykinesia (Bologna et al. 2023)). Consequently, models relying solely on a single task may yield suboptimal performance. To address this, we introduce a large video dataset featuring webcam recordings of individuals performing three tasks: (i) finger-tapping (motor function), (ii) smiling (facial expression), and (iii) uttering a pangram111pangram: sentence containing all the letters of the alphabet. (speech). The dataset is collected from 845 unique participants (272 with PD) from diverse demographics who recorded all of these tasks successfully (about 20% of them recorded multiple times), resulting in 1102 videos for each task and a total of 3306 videos. To the best of our knowledge, this is the first multi-task video dataset for PD screening.
We propose a novel two-stage classification model to distinguish between individuals with and without Parkinson’s disease (PD) using our dataset. First, each task is independently modeled with neural networks trained using Monte Carlo dropout (MC-dropout) to generate predictions and associated uncertainties. Next, we introduce the Uncertainty-calibrated Fusion Network (UFNet), which aggregates features from multiple tasks through an attention mechanism (Vaswani et al. 2017), while calibrating the attention scores based on task-specific uncertainties. UFNet is designed to produce PD/Non-PD predictions and uses MC-dropout during both training and inference to quantify prediction confidence. By filtering out low-confidence predictions, it increases patient safety. Our evaluations on a subject-separated test set demonstrate that UFNet significantly outperforms both single-task models and multi-modal fusion baselines. The model is computationally efficient, employing only a few shallow neural networks and a single self-attention module, making it suitable for deployment on smartphones or personal computers.
@@
  • We introduce the first-ever large-scale, multi-task video dataset for Parkinson’s disease screening collected from 845 individuals with diverse demography. Although we cannot share the raw videos to protect identifiable patient information, we will release the dataset in the form of extracted features (supplied as supplementary materials).
@@
Traditionally, PD is diagnosed by a clinician based on medical history and a clinical examination, which typically includes the completion of standardized tasks and rating each task according to the MDS-UPDRS guidance (Goetz et al. 2008). Recently, a cerebrospinal fluid (CSF) based $\alpha$-synuclein seed amplification assay has been developed (Siderowf et al. 2023), offering a potential diagnostic biomarker. However, access to clinical care can be limited, and diagnostic methods relying on the collection of CSF are invasive, costly, and burdensome to the patients.
Recent research has leveraged machine learning and sensors to facilitate remote assessment of PD. For example, nocturnal breathing signals, obtained from a breathing belt or reflected radio waves, can detect PD with high accuracy when analyzed by a machine learning model (Yang et al. 2022). Additionally, several body-worn sensors have been successfully used to monitor clinical features such as dyskinesia and gait disturbances associated with PD (Moreau et al. 2023). However, wearable sensors face challenges related to cost, comfort, and ease of use, which limit the scalability of these techniques for global adoption.
Using advanced computer vision and machine learning tools, precise clinical features can be extracted from recorded videos and used to screen for PD (Jin et al. 2020) or track symptom progression (Islam et al. 2023). However, existing methods for video analysis suffer from two major limitations – small cohort size (Jin et al. 2020), and reliance on single modality (Tsanas et al. 2009; Ali et al. 2020, 2021; Rahman et al. 2021; Islam et al. 2023; Adnan et al. 2023). Symptoms of PD are multi-faceted and may affect individuals differently. For instance, one individual may face speech difficulty but retain relatively normal motor functionality, while another individual may have prominent hypomimia (i.e., reduced facial expression) or bradykinesia (i.e., slowness incoordination of movements). Therefore, PD detection models may need to consider all these different modalities for improved efficacy.
This work addresses a critical gap in the field, which has persisted due to the lack of multimodal, diverse, and naturalistic data. Traditionally, clinical enterprises often lack the insight or foresight to determine what data to collect and in what quantities, while technical teams are constrained by limited access to patients and resources, preventing them from dedicating years to gathering such comprehensive data. These challenges have largely prevented the development of robust multimodal PD detection models. By establishing a new collaboration, we have been able to overcome these barriers and set new benchmarks for a critical disease. With the features we have made available, we hope the AI community will engage with this dataset and continue to advance the model further.
@@
We selected three standardized tasks that can be easily completed using a computer webcam and microphone, with or without external supervision:
  • (i)
    Finger-tapping: participants tap their thumb with the index finger ten times as fast as possible, first with the right hand, then with the left. Finger-tapping task is completed in accordance with the MDS-UPDRS scale to measure bradykinesia in the upper limb, a key sign of PD (Hughes et al. 1992).
  • (ii
    Smile: participants mimic a smile expression three times, alternating with a neutral face. While the expression may be unnatural, studies suggest it still captures signs of hypomimia (Bandini et al. 2017; Adnan et al. 2023).
  • (iii)
    Speech: participants utter a script, “The quick brown fox jumps over a lazy dog. The dog wakes up, and follows the fox into the forest. But again, the quick brown fox jumps over the lazy dog.” The first sentence is a pangram, containing all the letters of the English alphabet, and the later sentences are added to obtain a longer speech segment. Prior research identified this task as a promising way of screening PD (Rahman et al. 2021).
@@
We recruited a diverse cohort of participants (with and without PD) through multiple channels including a brain health study registry, social media, a PD wellness center, and clinician referrals. Among the approximately 1,400 individuals who recorded at least one task, 845 completed all three standardized tasks (comprising this dataset). Participants were primarily from the US, with diverse demographic characteristics: 52.7% female, a mean age of 61.9 years (22.5% over 70), and representation from various ethnic backgrounds. Among the 272 (32.2%) participants who had PD, 233 had their diagnosis confirmed through clinical evaluation, while 39 self-reported their condition. A detailed demographic summary of the participants is available in the supplementary materials (Technical Appendix A). This study was approved by relevant institutional review boards.
@@
Table 1: The single-task datasets contain videos of three different tasks: finger-tapping (motor function), smiling (facial expression), and speech (pangram utterance) from 1167, 1357, and 1265 participants, respectively. The multi-task dataset comprises 845 unique participants with videos of all three tasks. Validation and test set participants are the same across single-task and multi-task experiments (participants with missing videos are excluded).
@@
While most participants completed all three tasks, some did not complete one or more tasks, and some task videos were discarded due to feature extraction failures, often caused by inaccurate task completion. Instead of discarding videos of participants with missing tasks, we trained task-specific models with all available videos for each task. The multi-task model is then trained on the participants who completed all three tasks. We split the datasets based on the participants to ensure patient-centric evaluation. First, we listed all the participants ($n=1402$) enrolled in the study and then randomly assigned 60%, 20%, and 20% of the participants into the training, validation, and test sets, respectively. The random assignment was stratified to ensure a similar ratio of individuals with and without PD across these three folds. Both task-specific and multi-task models are validated and tested on the same participant cohort, with no model seeing any data from these participants during training. Table 1 provides a summary of the dataset splits.
@@
Figure 1: An illustrative overview of the proposed model. Task-specific models are shallow neural networks trained with task-specific features and MC-dropout. Standard deviations in multiple rounds of inference obtained from the task-specific models are used to calibrate the attention scores when fusing task-specific features. Finally, another shallow neural network is trained to differentiate individuals with and without PD, and withhold prediction when the model is uncertain. Authors have obtained consent to publish images of the human subject.
@@
We rely on prior literature to extract clinically meaningful features for each task. Although state-of-the-art deep learning models can learn to represent a video without explicit feature extraction (Liu et al. 2022b; Tong et al. 2022), training these models would require a notably larger dataset. By converting task-videos into feature sets, we can use simpler models with fewer trainable parameters.
@@
Islam et al. (2023) extracted 65 features to analyze the finger-tapping task for assessing PD severity. Using MediaPipe (Grishchenko et al. 2022) hand to detect hand movements and key-points, they measured clinically relevant features such as finger-tapping speed, amplitude, and interruptions. We apply this technique to both hands, extracting 130 features in total.
@@
We leveraged the 42 facial features extracted by Adnan et al. (2023) from smile mimicry videos. These features, captured using OpenFace (Baltrušaitis, Robinson, and Morency 2016) and MediaPipe, encompass key PD markers outlined by the MDS-UPDRS, such as eye blinking, lip separation, mouth opening, and intensity of facial muscle movements.
@@
We extracted 1024-dimensional embeddings from a pre-trained WavLM (Chen et al. 2022) language model to encode the pangram utterance task. WavLM, trained on massive amounts of speech data, excels at capturing acoustic characteristics of speech, making it useful for various tasks like speech recognition, speaker identification, and emotion recognition. Recent research has shown WavLM embeddings to be effective for PD screening (Adnan et al. 2024).
@@
  • Feature selection and scaling: Pairwise correlation among the features is calculated based on the training data. If two features have a Pearson’s correlation coefficient above a specified threshold, one feature is dropped. Feature values are scaled using StandardScaler or MinMaxScaler. Whether to apply feature selection or scaling, the correlation threshold, and the scaling method are hyper-parameters tuned on the validation set.
  • Shallow neural networks: The shallow neural network consists of 1-2 linear layers, with the last layer having a single output neuron. We use sigmoid activation with the output layer and ReLU with the other layers to add non-linearity.
  • Monte Carlo dropout (MC-dropout): To improve model robustness and estimate prediction uncertainty, we employ MC-dropout (Atighehchian et al. 2022). This technique allows using dropout for both training and inference. Therefore, we can obtain predictions from multiple rounds during testing. After $n$ rounds of inference, we obtain the mean ($\mu$) and standard deviation ($\sigma$) across $n$ different predictions for the same data. When the model is uncertain, we expect a relatively higher standard deviation across these predictions. Therefore, $\sigma$ is used to estimate model uncertainty.
The task-specific models are trained using binary cross-entropy loss, along with an optimizer (see Technical Appendix C regarding hyper-parameter tuning).
@@
Since the size of the features may vary from task to task, they are first projected to the same dimension ($d$) using a projection layer. Each projection layer consists of a linear layer ($\mathbb{R}^{d_{X_{i}}}\rightarrow\mathbb{R}^{d}$) with MC-dropout, followed by non-linear activation (ReLU) and layer normalization.
@@
We employ a self-attention mechanism to integrate the projected features ($X^{p}_{i}\in\mathbb{R}^{d}$) extracted from three distinct tasks. Unlike standard self-attention applications, our input sequence comprises task-specific feature vectors (sequence length = 3) rather than tokens of the same modality. For each task, projected features are transformed into query, key, and value representations using learned linear projections. Attention weights are computed based on the similarity between queries and keys.
@@
Here, $Q,K$ are the query and key matrices respectively, and $\eta$ is a hyper-parameter. After computing the attention scores, $X^{p}_{i}$ is converted into a contextualized representation $Z_{i}=\sum_{j=1}^{N}a^{\prime}_{j}v_{j}$, where $v_{j}$ is the corresponding value vector, and $N=3$ is the number of tasks. Please refer to Technical Appendix E for Pytorch implementation of the self-attention module.
@@
Using predictions obtained from multiple rounds of MC-dropout, we can model the confidence of the predictions. Here, we compute the $95\%$ confidence interval of the predicted scores, and if the interval contains the decision threshold (i.e., 0.50), the prediction is considered to be of low confidence. For patient safety, we withhold such predictions as they are more likely to be inaccurate.
Finally, the model is trained with binary cross-entropy loss and SGD optimizer with momentum $0.6898$. After hyper-parameter tuning on the validation set, the query dimension is set to 64 with a $0.0207$ learning rate, $0.4960$ dropout probability, and $\eta=81.8179$ (see Technical Appendix D for details).
@@
Experiments are conducted on an AMD Threadripper 3970x 32 core CPU with 256 GB RAM, and two NVIDIA A6000 GPU with 48 GB VRAM. We use Python (PyTorch deep learning framework) and Linux operating system.
@@
Majority Voting. The predictions from three different task-specific models are combined to generate a single prediction. The final prediction is the class (PD or Non-PD) agreed upon by the majority (i.e., two or more).
@@
Hybrid Fusion Baseline. Task-specific features and prediction scores (logits) are both provided as input to a shallow neural network. This network leverages both the input features and prediction scores, combining the strengths of early and late fusion approaches.
Ablation Models. UFNet uses task-specific features and prediction scores, making it a hybrid fusion approach. Additionally, we analyze the effect of removing task-specific predictions from UFNet, implementing an early fusion approach. We also examine the impact of withholding predictions for both early and hybrid versions of the UFNet model.
Model Selection and Performance Reporting
The model with the highest AUROC on the validation set is selected for the reported experiments. After selection, each model is run with 30 different random seeds, and evaluated on the test set. The average and 95% confidence interval of each performance metric are reported. Due to dataset imbalance, we report accuracy, balanced accuracy, $F_{1}$ score, AUROC, and AUPRC to compare the models’ performance. Specificity, sensitivity, positive predictive value (PPV), and negative predictive value (NPV) of UFNet and other baselines are also reported. Coverage (of cases with a prediction) is provided when uncertain predictions are withheld. Significant differences mean no overlap in 95% confidence intervals; non-inferiority means overlap.
For patient safety, it is critical to ensure that the prediction scores approximate the true probability of the positive class. Therefore, we evaluate model calibration using expected calibration error (ECE) (Nixon et al. 2019) and Brier score (Rufibach 2010). Misclassification rates across male and female subgroups are compared using a two-sample $Z$-test for proportions to analyze potential bias. Model performance across white and non-white participants is compared using Fisher’s exact test due to the small number of non-white participants available in the test set. We also report average misclassification rates and confidence intervals across different age groups.

### v4

@@ Abstract
Limited accessibility to neurological care leads to under-diagnosed Parkinson’s Disease (PD), preventing early intervention. Existing AI-based PD detection methods primarily focus on unimodal analysis of motor or speech tasks, overlooking the multifaceted nature of the disease. To address this, we introduce a large-scale, multi-task video dataset of $1102$ sessions (each containing videos of finger tapping, facial expression, and speech tasks captured via webcam) from $845$ participants ($272$ with PD). We propose a novel Uncertainty-calibrated Fusion Network (UFNet) that leverages this multimodal data to enhance diagnostic accuracy. UFNet employs independent task-specific networks, trained with Monte Carlo Dropout for uncertainty quantification, followed by self-attended fusion of features, with attention weights dynamically adjusted based on task-specific uncertainties. We randomly split the participants into training (60%), validation (20%), and test (20%) sets to ensure patient-centered evaluation. UFNet significantly outperformed single-task models in terms of accuracy, area under the ROC curve (AUROC), and sensitivity while maintaining non-inferior specificity. Withholding uncertain predictions further boosted the performance, achieving $88.0\pm 0.3\%$ accuracy, $93.0\pm 0.2\%$ AUROC, $79.3\pm 0.9\%$ sensitivity, and $92.6\pm 0.3\%$ specificity, at the expense of not being able to predict for $2.3\pm 0.3\%$ data ($\pm$ denotes 95% confidence interval). Further analysis suggests that the trained model does not exhibit any detectable bias across sex and ethnic subgroups and is most effective for individuals aged between 50 and 80. By merely requiring a webcam and microphone, our approach facilitates accessible home-based PD screening, especially in regions with limited healthcare resources.
Code & DatasetTBA
Published versionTBA
Demohttps://parktest.net/demo
@@ Introduction
Detecting PD is particularly challenging due to the variability of individual symptoms. For example, while vocal impairment is a common PD symptom (Ho et al. 1999), patients may also exhibit PD through other modalities, such as facial expression (e.g., hypomimia (Gunnery et al. 2016)) or motor function (e.g., bradykinesia (Bologna et al. 2023)). Consequently, models relying solely on a single task may yield suboptimal performance. To address this, we introduce a large video dataset featuring webcam recordings of individuals performing three tasks: (i) finger-tapping (motor function), (ii) smiling (facial expression), and (iii) uttering a pangram111pangram: sentence containing all the letters of the alphabet. (speech). The dataset is collected from 845 unique participants (272 with PD) from diverse demographics who recorded all of these tasks successfully (about 20% of them recorded multiple times), resulting in 1102 videos for each task and a total of 3306 videos. To our knowledge, this is the first multi-task video dataset for PD screening.
We propose a novel two-stage classification model to distinguish between individuals with and without Parkinson’s disease (PD) using our dataset. First, each task is independently modeled with neural networks trained using Monte Carlo dropout (MC dropout) to generate predictions and associated uncertainties. Next, we introduce the Uncertainty-calibrated Fusion Network (UFNet), which aggregates features from multiple tasks through an attention mechanism (Vaswani et al. 2017), while calibrating the attention scores based on task-specific uncertainties. UFNet is designed to produce PD/Non-PD predictions and uses MC dropout during training and inference to quantify prediction confidence. By filtering out low-confidence predictions, it increases patient safety. Our evaluations on a subject-separated test set demonstrate that UFNet significantly outperforms single-task models and multi-modal fusion baselines. The model is computationally efficient, employing only a few shallow neural networks and a single self-attention module, making it suitable for deployment on smartphones or personal computers.
@@
  • We introduce the first-ever large-scale, multi-task video dataset for Parkinson’s disease screening collected from 845 individuals with diverse demography. Although we cannot share the raw videos to protect identifiable patient information, we release the dataset as extracted features.
@@
Traditionally, Parkinson’s Disease (PD) is diagnosed by a clinician based on medical history and a clinical examination, typically including completing standardized tasks and rating each task according to the MDS-UPDRS guidance (Goetz et al. 2008). Recently, a cerebrospinal fluid (CSF) based $\alpha$-synuclein seed amplification assay has been developed (Siderowf et al. 2023), offering a potential diagnostic biomarker. However, access to clinical care can be limited, and diagnostic methods relying on the collection of CSF are invasive, costly, and burdensome to the patients.
Recent research has leveraged machine learning (ML) and sensors to enable remote assessment of PD. For example, breathing signals obtained from reflected radio waves can detect PD accurately when analyzed by an ML model (Yang et al. 2022). Additionally, body-worn sensors have been successfully used to monitor clinical features such as dyskinesia and gait disturbances associated with PD (Moreau et al. 2023). However, wearable sensors face challenges related to cost, comfort, and ease of use, which limit the scalability of these techniques for global adoption.
Advanced pose tracking models like MediaPipe (Lugaresi et al. 2019) enable extracting precise clinical features from recorded videos, later used for PD screening (Jin et al. 2020) or progression tracking (Islam et al. 2023). However, existing methods for video analysis suffer from major limitations – small cohort size (Jin et al. 2020), and reliance on single modality (Tsanas et al. 2009; Ali et al. 2020, 2021; Rahman et al. 2021; Adnan et al. 2023). Symptoms of PD are multi-faceted and may affect individuals differently. For instance, one individual may face speech difficulty but retain normal motor functionality, while others may have prominent hypomimia (i.e., reduced facial expression) or bradykinesia (i.e., slowness of movement). Therefore, PD detection models may need to consider multiple modalities for improved efficacy.
This work addresses a critical gap in the field, which has persisted due to the lack of multimodal, diverse, and naturalistic data. By utilizing the PARK framework (Langevin et al. 2019) and inter-disciplinary collaboration, we have collected videos of multiple tasks from a large, diverse cohort, enabling us to develop an accurate PD screening model. We hope the community will engage with this dataset and continue to advance the model.
@@
We select three standardized tasks that can be easily completed using a computer webcam and microphone, with or without external supervision:
Finger-tapping: Participants tap their thumb with the index finger ten times as fast as possible, first with the right hand, then with the left. The finger-tapping task is completed following the MDS-UPDRS scale to measure bradykinesia in the upper limb, a key sign of PD (Hughes et al. 1992).
Smile: Participants mimic a smile expression three times, alternating with a neutral face. Studies suggest the task captures signs of hypomimia (Bandini et al. 2017; Adnan et al. 2023), even though the expression may be unnatural.
Speech: Participants utter a script, “The quick brown fox jumps over a lazy dog. The dog wakes up and follows the fox into the forest. But again, the quick brown fox jumps over the lazy dog.” The first sentence is a pangram, containing all the letters of the English alphabet, and the later sentences are added to obtain a longer speech segment. Prior research identified this task as a promising way of screening PD (Rahman et al. 2021).
@@
We recruited participants (with and without PD) from a diverse background (see Table 1) through multiple channels, including a brain health study registry, social media, a PD wellness center, and clinician referrals. Among the approximately 1,400 individuals who recorded at least one task, 845 completed all three standardized tasks (comprising this dataset). Participants are primarily from the US, with diverse demographic characteristics: 52.7% female, a mean age of 61.9 years (22.5% over 70), and representation from various ethnic backgrounds. Among the 272 (32.2%) participants who had PD, 233 had their diagnosis confirmed through clinical evaluation, while 39 self-reported their condition. This study was approved by the institutional review board (IRB) of the University of Rochester.
Subgroup Attribute With PD Without PD Total
Number of participants 272 573 845
Sex, n(%) Female 122 (44.9%) 323 (56.4 %) 445 (52.7%)
Male 147 (54.0%) 250 (43.6 %) 397 (47.0%)
Nonbinary 1 (0.4%) 0 (0.0 %) 1 (0.1%)
Unknown 2 (0.7%) 0 (0.0 %) 2 (0.2%)
Age in years, n (%)
(range: 18.0 - 93.0, mean: 61.9)
Below 20 0 (0.0 %) 6 (1.0 %) 6 (0.7 %)
20-29 1 (0.4 %) 28 (4.9 %) 29 (3.4 %)
30-39 2 (0.7 %) 19 (3.3 %) 21 (2.5 %)
40-49 6 (2.2 %) 17 (3.0 %) 23 (2.7 %)
50-59 33 (12.1 %) 119 (20.8 %) 152 (18.0 %)
60-69 94 (34.6 %) 231 (40.3 %) 325 (38.5 %)
70-79 98 (36.0 %) 76 (13.3 %) 174 (20.6 %)
80 and above 12 (4.4 %) 4 (0.7 %) 16 (1.9 %)
Unknown 26 (9.6 %) 73 (12.7 %) 99 (11.7 %)
Ethnicity, n (%) American Indian or Alaska Native 1 (0.4 %) 0 (0.0 %) 1 (0.1%)
Asian 3 (1.1 %) 34 (5.9 %) 37 (4.4%)
Black or African American 3 (1.1 %) 29 (5.1 %) 32 (3.8%)
white 163 (59.9 %) 463 (80.8 %) 626 (74.1%)
Others 2 (0.7 %) 3 (0.5 %) 5 (0.6%)
Unknown 100 (36.8 %) 44 (7.7 %) 144 (17.0%)
Disease duration in years, n (%)
(range: 1.0 - 24.0, mean: 6.6)
<=2 19 (6.99%) - -
2-5 36 (13.24%) - -
5-10 28 (10.29%) - -
10-15 10 (3.68%) - -
15-20 7 (2.57%) - -
>20 2 (0.74%) - -
Unknown 170 (62.5%) - -
Cohort, n (%) Home 39 (14.3 %) 399 (69.6 %) 438 (51.8 %)
Clinic 91 (33.5 %) 107 (18.7 %) 198 (23.4 %)
PD wellness center 142 (52.2 %) 67 (11.7 %) 209 (24.7 %)
Table 1: Summary of demographic information. Participants recruited via clinician referrals (Clinic cohort) and the PD wellness center were clinically diagnosed, but the other participants’ diagnosis status was self-reported. Information related to sex, age, ethnicity, and disease duration were self-reported by the participants. It is important to note that demographic information was optional, and missing data is indicated as “Unknown.”
@@
Table 2: Summary of dataset splits.
@@
Most participants completed all three tasks, but some did not, and some task videos were discarded due to feature extraction failures, often caused by inaccurate task completion. Instead of discarding videos of participants with missing tasks, we trained task-specific models with all available videos for each task. The multi-task model is then trained on the participants who completed all three tasks (Table 2). We split the datasets based on the participants to ensure patient-centric evaluation. All participants were randomly assigned into training (60%), validation (20%), and test (20%) sets, with stratification to maintain a similar ratio of individuals with and without PD across folds. Both task-specific and multi-task models are validated and tested on the same participant cohort, with no model seeing any data from these participants during training.
@@
Figure 1: An illustrative overview of the proposed model. Task-specific models are shallow neural networks trained with task-specific features and MC dropout. Standard deviations in multiple rounds of inference obtained from the task-specific models are used to calibrate the attention scores when fusing task-specific features. Finally, another shallow neural network is trained to differentiate individuals with and without PD and withhold prediction when the model is uncertain.
@@
We rely on prior literature to extract task-specific video features. For the smile and finger-tapping tasks, clinically meaningful hand-crafted features were the better choice against deep video models such as ViViT (Arnab et al. 2021), TimeSformer (Bertasius, Wang, and Torresani 2021), VideoMAE (Tong et al. 2022), and Uniformer (Li et al. 2023), based on their performance and suitability for low-resource deployment. WavLM (Chen et al. 2022) has been identified as the most effective for speech-based PD screening by a recent work (Adnan et al. 2024).
@@
Islam et al. (2023) extracted 65 features to analyze the finger-tapping task for assessing PD severity. Using MediaPipe hand (Grishchenko et al. 2022) to detect hand movements and key points, they measured clinically relevant features such as finger-tapping speed, amplitude, and interruptions. We apply this technique to both hands, extracting 130 features in total.
@@
We used 42 facial features from smile videos extracted by  Adnan et al. (2023). These features, extracted using OpenFace (Baltrušaitis, Robinson, and Morency 2016) and MediaPipe, capture key PD markers such as eye blinking, lip separation, mouth opening, and intensity of facial muscle movements.
@@
We extracted 1024-dimensional embeddings from a pre-trained WavLM (Chen et al. 2022) model to encode the speech task. WavLM, trained on massive amounts of speech data, excels at capturing acoustic characteristics of speech, making it useful for tasks like speaker identification, speech, and emotion recognition.
@@
  • Feature selection and scaling: We calculate pairwise correlation among the features to drop highly correlated features. Feature values are then scaled using StandardScaler or MinMaxScaler. Whether to apply feature selection or scaling, the correlation threshold and the scaling method are hyper-parameters tuned on the validation set.
  • Shallow neural networks: The network consists of 1-2 linear layers, with a single output neuron in the last layer. We use sigmoid activation with the output layer and ReLU with the other layers to add non-linearity.
  • Monte Carlo dropout (MC dropout): To improve model robustness and estimate prediction uncertainty, we employ MC dropout (Atighehchian et al. 2022). This technique allows using dropout for both training and inference. Therefore, we can obtain predictions from multiple rounds during testing and the mean ($\mu$) and standard deviation ($\sigma$) of these predictions. We expect a relatively higher standard deviation across these predictions when the model is uncertain. Therefore, $\sigma$ is used to estimate model uncertainty.
The task-specific models are trained using binary cross-entropy loss and either SGD or AdamW optimizer.
@@
Since the size of the features may vary from task to task, they are first projected to the same dimension ($d$) using a projection layer. Each projection layer consists of a linear layer ($\mathbb{R}^{d_{X_{i}}}\rightarrow\mathbb{R}^{d}$) with MC dropout, followed by non-linear activation (ReLU) and layer normalization.
@@
We employ self-attention to integrate the projected features ($X^{p}_{i}\in\mathbb{R}^{d}$) extracted from three distinct tasks. Unlike standard self-attention applications, our input sequence comprises task-specific feature vectors (sequence length = 3) rather than tokens of the same modality. For each task, projected features are transformed into query ($Q$), key ($K$), and value ($V$) vectors using linear projections. Attention weights are computed based on the similarity between queries and keys.
@@
Hyper-parameter $\eta$ controls the degree of uncertainty-based down-weighting. After computing the attention scores, $X^{p}_{i}$ is converted into a contextualized representation $Z_{i}=\sum_{j=1}^{N}A_{i,j}V_{j}$, where $N=3$ is the number of tasks.
@@
Our proposed approach is designed for neurological screening, specifically to suggest clinical evaluation for individuals if the model identifies PD symptoms. Since mispredictions in this context could lead to significant consequences (e.g., unnecessary anxiety or delayed care), the model is designed to withhold predictions it identifies as uncertain. To achieve this, we implement two different techniques:
  • MC Dropout: For new test data, we generate $n$ predictions by running multiple rounds of Monte Carlo (MC) dropout. We compute the $95\%$ confidence interval (CI) for these predictions and classify them as uncertain if the CI contains the decision threshold (i.e., 0.50).
  • Conformal Prediction: We partition the training data into a training set (80%) and a calibration set (20%). We calculate a conformity threshold using the calibration set and identify uncertain predictions when the prediction set contains both PD and Non-PD classes.
We also experiment with performing label smoothing (Müller, Kornblith, and Hinton 2019) (which redistributes probability mass to reduce overconfidence) and/or Platt scaling (Platt et al. 1999) (which fits a logistic regression model to improve calibration) with conformal prediction (Saunders, Gammerman, and Vovk 1999) to see whether they further improve model performance.
Hyper-parameter tuning.
We utilized Weights & Biases [^2] https://wandb.ai to employ Bayesian hyper-parameter tuning within the defined search space (see Appendix B and C) to maximize the AUROC on the validation set. Finally, the UFNet model is trained with binary cross-entropy loss and SGD optimizer with learning rate $0.0207$ and momentum $0.6898$. We selected 64 as the dimension for the query, key, and value vectors. The chosen value of dropout probability was $0.4960$, and $\eta$ was 81.8179. Since some hyper-parameters had continuous search space (e.g., the learning rate was tuned within $1e^{-5}$ and 1.0), we report the precise values of the selected hyper-parameters.
@@
Experiments are conducted on an AMD Threadripper 3970x 32-core CPU with 256 GB RAM and two NVIDIA A6000 GPUs with 48 GB VRAM. We use Python (PyTorch deep learning framework) and Linux operating system.
@@
Majority Voting. The three task-specific models’ predictions are combined to generate a single prediction. The final prediction is the class (PD or Non-PD) agreed upon by the majority (i.e., two or more).
@@
Hybrid Fusion Baseline. Task-specific features and prediction scores (logits) are provided as input to a shallow neural network. This network leverages both the input features and prediction scores, combining the strengths of early and late fusion approaches.
Performance Reporting
After hyper-parameter tuning, we train the selected model with 30 random seeds and evaluate it on the test set. We report the average and 95% confidence interval (CI) of each performance metric. In addition to standard binary classification metrics, we report coverage (of cases where the model is confident enough to make a prediction) when appropriate. Furthermore, ensuring that the class probabilities output by a model approximates the actual probability is critical for patient safety. Therefore, we evaluate model calibration using expected calibration error (ECE) (Nixon et al. 2019) and Brier score (Rufibach 2010). Unless specified otherwise, a significant difference is identified by having no overlap in the 95% CI, while non-inferiority indicates overlap.